OpenStudy (rockinhood):

+MEDAL WILL BE AWARDED!+ Carl deposited $5,142 into a savings account 20 years ago. The account has an interest rate of 3.8% and the balance is currently $10,994.58. How often does the interest compound? daily monthly quarterly annually

OpenStudy (rockinhood):

@sweetburger ? Perhaps @Nnesha ?

OpenStudy (mertsj):

\[A=P(1+\frac{r}{n})^{nt}\] \[10994.58=5142(1+\frac{.038}{n})^{20n}\]

OpenStudy (rockinhood):

5,142(1 + .038/365)^7300 I found was = to what I needed

OpenStudy (rockinhood):

So daily is my answer? http://www.wolframalpha.com/input/?i=5%2C142%281+%2B+.038%2F365%29%5E7300